# Explaining Object Detectors via Collective Contribution of Pixels

### v2

@@ D.3 Sanity checks
We perform the checks using model weight randomization. Figure J shows an example generated by VX-CODE with patch insertion for DETR. The proposed method identifies uninformative patches under randomized weights. Table 5 presents the SSIM and rank correlation results computed for 100 detected instances using DETR on MS-COCO. These metrics evaluate the similarity between the explanations for the original and randomized models (with values closer to 1.0 indicating higher similarity). We observe significant drops in these metrics. These results indicate that the proposed method provides meaningful explanations that are sensitive to the model’s parameters.
@@ D.5 Additional results for visualizations
In Sec. 5.2, we show the explanations generated by each method. Figure K and Fig. L provide the additional examples of explanations generated by each method for detections from DETR and Faster R-CNN, respectively. Each figure includes the heat maps generated by each method and the patches identified by VX-CODE in the early steps. As described in Sec. 5.2, the proposed method identifies the important patches for detections by considering the collective contributions of multiple patches.
@@ D.6 Additional results for patch combination
In Sec. 5.2, we demonstrate the effect of increasing $r$. Here, we show the additional results to further support this analysis. Figure M illustrates the effect of $r$ on patch identification. As shown in Fig. M, the proposed method identifies more features, reflecting the effect of considering the collective contributions of a larger set of patches. For example, as shown on the left side of Fig. M (a) (detection for the aeroplane), the identified features for $r=1$ include the front and rear parts of the aeroplane, while the wings are additionally identified for $r=2$ and $r=3$, indicating that these features collectively contribute to the detection.
Increasing $r$ also leads to more accurate identification of important patches. As shown on the right side of Fig. M (a) (detection of the banana), the case with $r=1$ fails to identify the critical patches in the initial steps, whereas $r=2$ and $r=3$ successfully identify these patches at earlier steps. This improvement is particularly effective for patch insertion. Patch insertion gradually adds patches to an empty image. Therefore, at earlier steps, the reward function value changes only slightly due to the limited pixel information, making it difficult to identify important patches. Increasing $r$ mitigates this problem by considering Shapley values and interactions among more patches, leading to more accurate identification of important patches.
Figure O: Additional results comparing the visualizations generated by each method for the biased model. In VX-CODE, we set $r=1$. The results correspond to the classes aeroplane, bird, bottle, bus, dog, motorbike, and sofa, listed from the top row downward.
Figure P: Patches identified from VX-CODE with patch insertion (left) and patch deletion (right) corresponding to the heat maps shown in Fig. O.
@@ D.7 Additional results for biased model
Figure N illustrates an example of the constructed synthetic dataset. As shown in Fig. N (a), a square marker is placed outside the upper-right corner of the bounding box for seven classes: aeroplane, bird, bottle, bus, dog, motorbike, and sofa. These data include annotations for the corresponding class and its bounding box. In our constructed dataset, these synthetic data account for approximately 70%. In the remaining 30%, data corresponding to each class but without annotations are included and are treated as background during training. DETR trained on such data considers both the features of the marker outside the bounding box and those of the instance itself to detect objects and distinguish classes appropriately. In practice, this model achieves a mean average precision (mAP@50) with an IoU threshold >0.5>0.5 of 83.41 for the seven classes on the synthetic test dataset. On the other hand, the mAP@50 drops to 34.73 on the original test dataset, where the square marker is not placed. These results indicate that DETR trained on the constructed synthetic dataset is heavily biased toward the square marker outside the bounding box.
@@
Figure O shows additional examples of visualizations generated by each method for the biased model, and Fig. P shows the patches identified from VX-CODE. As discussed in Sec 4, the proposed method efficiently identifies the features of both the instance and the marker, whereas existing methods tend to highlight only the marker (e.g., the result of Grad-CAM++ and D-RISE for the detection of the dog) or emphasize all features (e.g., the result of SSGrad-CAM++ for the detection of the bottle).
Figure Q: Comparison results of the visual explanations for bounding box predictions generated by each method. In VX-CODE, we set $r=1$. (a) Heat maps generated by each method. (b) Patches identified from VX-CODE with patch insertion (left) and patch deletion (right). Each heat map corresponds to those shown in (a).
Figure R: The differences in regions identified by VX-CODE when varying $\alpha$ in Eq. (129), with (a) patch insertion and (b) patch deletion.
Figure S: The differences in identified regions by VX-CODE with patch insertion between (a) DETR and (b) Faster R-CNN. Each figure shows the number of identified patches $k$ and the corresponding reward value.
@@ D.8 Quantitative evaluation for class explanations
Figure T: Visualization results for misclassification generated by D-RISE, SSGrad-CAM, ODAM, SSGrad-CAM++, and VX-CODE with patch insertion. In the first row, the horse is misclassified as a bird; in the second row, the bus is misclassified as a train; and in the third row, the horse is misclassified as a dog.
Figure U: Visualization results for mislocalization generated by D-RISE, SSGrad-CAM, ODAM, SSGrad-CAM++, and VX-CODE with patch insertion. In each detection result, the green box represents the ground truth bounding box, while the red box represents the predicted bounding box. The predicted labels, from top to bottom, are person, bicycle, and chair.
@@ D.9 Analysis for bounding box explanations
Table 10 shows the results of insertion, deletion, and overall metrics, based on 1,000 instances from the results of DETR on MS-COCO. Here, we use Eq. (127) as the similarity score calculated at each step. The proposed method outperforms the others, demonstrating its ability to accurately identify the important regions for the bounding box generation. Figure Q compares the visualization results. Generally, the size of a bounding box is determined by several points (at least two). Therefore, considering their collective contributions is important. As shown in Fig. Q, the proposed method accurately identifies important regions by considering such collective contributions. In contrast, D-RISE, which uses the same reward function, and methods utilizing gradients generate noisy results or also highlight irrelevant regions. These results support its superiority in the quantitative results shown in Tab. 10.
@@ D.10 Multi-perspective explanations for detection results
Figure R shows the identified patches and the generated heat maps from VX-CODE with (a) patch insertion and (b) patch deletion in the cases where $\alpha$ is $0.01$, $0.50$, and $0.99$ These examples illustrate how the focus shifts between bounding box prediction and class prediction depending on the weight in the reward function. For example, as shown in the left example of (a) patch insertion, for the detection of the surfboard with $\alpha=0.01$, the proposed method preferentially highlights area near the front and rear of the surfboard, indicating that these features significantly contribute to the bounding box prediction. In the case with $\alpha=0.99$, the proposed method preferentially highlights the rear of the surfboard and the sea outside the bounding box, indicating that these features significantly contribute to the class (surfboard) prediction. In the case with $\alpha=0.50$, the proposed method highlights these features (the front and rear of the surfboard and the sea outside the bounding box) in a balanced manner. A similar trend can be seen in the examples of (b) patch deletion. For example, as shown in the left example, for the detection of the apple with $\alpha=0.01$, the proposed method preferentially highlights regions near the bounding box outline. In the case with $\alpha=0.99$, the proposed method not only highlights the detected apple but also other apples. This identification is natural because the selection focuses on the class prediction of apple, and these features influence it in other proposals generated by the object detector. In the case with $\alpha=0.50$, the proposed method only highlights features of the apple within the detected instance. As shown in these examples, the proposed method can adjust its focus between bounding box generation and class prediction by adjusting the weight in the reward function, enabling the generation of multi-perspective explanations.
@@ D.11 Analysis of object detectors using VX-CODE
Figure S shows the identified patches and generated heat maps for detections of a banana produced by DETR and Faster R-CNN. For DETR, patches in multiple parts are identified in the early steps, and the reward value defined in Eq. (3) reaches $0.91$ after four patches are identified. In contrast, for Faster R-CNN, patches in specific regions are preferentially identified, and the reward value reaches $0.94$ after 14 patches are identified. These results indicate that DETR, a transformer-based architecture, recognizes instances in a more compositional manner and with less information compared to the CNN-based architecture Faster R-CNN.
@@ D.12 Additional results for failure cases
In Secs. 4 and 5.3, we discuss explanations for failure cases. Here, we provide additional results. Figure T and Fig. U show the explanations generated by each method for misclassification and mislocalization, respectively. These results demonstrate the effectiveness of the proposed method in analyzing failure cases. For example, as shown in the first row of Fig. T, where the object detector misclassifies the horse as a bird, the proposed method identifies not only the animal’s body but also the tree outside the bounding box, indicating that these features (the animal’s body and the tree) contribute to the misdetection of the bird. In contrast, other methods do not highlight the tree outside the bounding box and fail to capture the effect of these features. The proposed method is also effective in analyzing mislocalization. For example, as shown in the first row of Fig. U, where the object detector mislocalizes the person by including the chair, the proposed method identifies both the features of the person and the towel covering the chair, indicating that these features contributed to the mislocalization. In contrast, other methods either highlight only one of these features or produce noisy explanations.

### v4

@@ D.3 Sanity checks
We perform the checks using model weight randomization. Figure K shows an example generated by VX-CODE with patch insertion for DETR. The proposed method identifies uninformative patches under randomized weights. Table 5 presents the SSIM and rank correlation results computed for 100 detected instances using DETR on MS-COCO. These metrics evaluate the similarity between the explanations for the original and randomized models (with values closer to 1.0 indicating higher similarity). We observe significant drops in these metrics. These results indicate that the proposed method provides meaningful explanations that are sensitive to the model’s parameters.
@@ D.5 Additional results for visualizations
In Sec. 5.2, we show the explanations generated by each method. Figure L and Fig. M provide the additional examples of explanations generated by each method for detections from DETR and Faster R-CNN, respectively. Each figure includes the heat maps generated by each method and the patches identified by VX-CODE in the early steps. As described in Sec. 5.2, the proposed method identifies the important patches for detections by considering the collective contributions of multiple patches.
@@ D.6 Additional results for patch combination
In Sec. 5.2, we demonstrate the effect of increasing $r$. Here, we show the additional results to further support this analysis. Figure N illustrates the effect of $r$ on patch identification. As shown in Fig. N, the proposed method identifies more features, reflecting the effect of considering the collective contributions of a larger set of patches. For example, as shown on the left side of Fig. N (a) (detection for the aeroplane), the identified features for $r=1$ include the front and rear parts of the aeroplane, while the wings are additionally identified for $r=2$ and $r=3$, indicating that these features collectively contribute to the detection.
Increasing $r$ also leads to more accurate identification of important patches. As shown on the right side of Fig. N (a) (detection of the banana), the case with $r=1$ fails to identify the critical patches in the initial steps, whereas $r=2$ and $r=3$ successfully identify these patches at earlier steps. This improvement is particularly effective for patch insertion. Patch insertion gradually adds patches to an empty image. Therefore, at earlier steps, the reward function value changes only slightly due to the limited pixel information, making it difficult to identify important patches. Increasing $r$ mitigates this problem by considering Shapley values and interactions among more patches, leading to more accurate identification of important patches.
Figure P: Additional results comparing the visualizations generated by each method for the biased model. In VX-CODE, we set $r=1$. The results correspond to the classes aeroplane, bird, bottle, bus, dog, motorbike, and sofa, listed from the top row downward.
Figure Q: Patches identified from VX-CODE with patch insertion (left) and patch deletion (right) corresponding to the heat maps shown in Fig. P.
@@ D.7 Additional results for biased model
Figure O illustrates an example of the constructed synthetic dataset. As shown in Fig. O (a), a square marker is placed outside the upper-right corner of the bounding box for seven classes: aeroplane, bird, bottle, bus, dog, motorbike, and sofa. These data include annotations for the corresponding class and its bounding box. In our constructed dataset, these synthetic data account for approximately 70%. In the remaining 30%, data corresponding to each class but without annotations are included and are treated as background during training. DETR trained on such data considers both the features of the marker outside the bounding box and those of the instance itself to detect objects and distinguish classes appropriately. In practice, this model achieves a mean average precision (mAP@50) with an IoU threshold >0.5>0.5 of 83.41 for the seven classes on the synthetic test dataset. On the other hand, the mAP@50 drops to 34.73 on the original test dataset, where the square marker is not placed. These results indicate that DETR trained on the constructed synthetic dataset is heavily biased toward the square marker outside the bounding box.
@@
Figure P shows additional examples of visualizations generated by each method for the biased model, and Fig. Q shows the patches identified from VX-CODE. As discussed in Sec 4, the proposed method efficiently identifies the features of both the instance and the marker, whereas existing methods tend to highlight only the marker (e.g., the result of Grad-CAM++ and D-RISE for the detection of the dog) or emphasize all features (e.g., the result of SSGrad-CAM++ for the detection of the bottle).
Figure R: Comparison results of the visual explanations for bounding box predictions generated by each method. In VX-CODE, we set $r=1$. (a) Heat maps generated by each method. (b) Patches identified from VX-CODE with patch insertion (left) and patch deletion (right). Each heat map corresponds to those shown in (a).
Figure S: The differences in regions identified by VX-CODE when varying $\alpha$ in Eq. (129), with (a) patch insertion and (b) patch deletion.
Figure T: The differences in identified regions by VX-CODE with patch insertion between (a) DETR and (b) Faster R-CNN. Each figure shows the number of identified patches $k$ and the corresponding reward value.
@@ D.8 Quantitative evaluation for class explanations
Figure U: Visualization results for misclassification generated by D-RISE, SSGrad-CAM, ODAM, SSGrad-CAM++, and VX-CODE with patch insertion. In the first row, the horse is misclassified as a bird; in the second row, the bus is misclassified as a train; and in the third row, the horse is misclassified as a dog.
Figure V: Visualization results for mislocalization generated by D-RISE, SSGrad-CAM, ODAM, SSGrad-CAM++, and VX-CODE with patch insertion. In each detection result, the green box represents the ground truth bounding box, while the red box represents the predicted bounding box. The predicted labels, from top to bottom, are person, bicycle, and chair.
@@ D.9 Analysis for bounding box explanations
Table 10 shows the results of insertion, deletion, and overall metrics, based on 1,000 instances from the results of DETR on MS-COCO. Here, we use Eq. (127) as the similarity score calculated at each step. The proposed method outperforms the others, demonstrating its ability to accurately identify the important regions for the bounding box generation. Figure R compares the visualization results. Generally, the size of a bounding box is determined by several points (at least two). Therefore, considering their collective contributions is important. As shown in Fig. R, the proposed method accurately identifies important regions by considering such collective contributions. In contrast, D-RISE, which uses the same reward function, and methods utilizing gradients generate noisy results or also highlight irrelevant regions. These results support its superiority in the quantitative results shown in Tab. 10.
@@ D.10 Multi-perspective explanations for detection results
Figure S shows the identified patches and the generated heat maps from VX-CODE with (a) patch insertion and (b) patch deletion in the cases where $\alpha$ is $0.01$, $0.50$, and $0.99$ These examples illustrate how the focus shifts between bounding box prediction and class prediction depending on the weight in the reward function. For example, as shown in the left example of (a) patch insertion, for the detection of the surfboard with $\alpha=0.01$, the proposed method preferentially highlights area near the front and rear of the surfboard, indicating that these features significantly contribute to the bounding box prediction. In the case with $\alpha=0.99$, the proposed method preferentially highlights the rear of the surfboard and the sea outside the bounding box, indicating that these features significantly contribute to the class (surfboard) prediction. In the case with $\alpha=0.50$, the proposed method highlights these features (the front and rear of the surfboard and the sea outside the bounding box) in a balanced manner. A similar trend can be seen in the examples of (b) patch deletion. For example, as shown in the left example, for the detection of the apple with $\alpha=0.01$, the proposed method preferentially highlights regions near the bounding box outline. In the case with $\alpha=0.99$, the proposed method not only highlights the detected apple but also other apples. This identification is natural because the selection focuses on the class prediction of apple, and these features influence it in other proposals generated by the object detector. In the case with $\alpha=0.50$, the proposed method only highlights features of the apple within the detected instance. As shown in these examples, the proposed method can adjust its focus between bounding box generation and class prediction by adjusting the weight in the reward function, enabling the generation of multi-perspective explanations.
@@ D.11 Analysis of object detectors using VX-CODE
Figure T shows the identified patches and generated heat maps for detections of a banana produced by DETR and Faster R-CNN. For DETR, patches in multiple parts are identified in the early steps, and the reward value defined in Eq. (3) reaches $0.91$ after four patches are identified. In contrast, for Faster R-CNN, patches in specific regions are preferentially identified, and the reward value reaches $0.94$ after 14 patches are identified. These results indicate that DETR, a transformer-based architecture, recognizes instances in a more compositional manner and with less information compared to the CNN-based architecture Faster R-CNN.
@@ D.12 Additional results for failure cases
In Secs. 4 and 5.3, we discuss explanations for failure cases. Here, we provide additional results. Figure U and Fig. V show the explanations generated by each method for misclassification and mislocalization, respectively. These results demonstrate the effectiveness of the proposed method in analyzing failure cases. For example, as shown in the first row of Fig. U, where the object detector misclassifies the horse as a bird, the proposed method identifies not only the animal’s body but also the tree outside the bounding box, indicating that these features (the animal’s body and the tree) contribute to the misdetection of the bird. In contrast, other methods do not highlight the tree outside the bounding box and fail to capture the effect of these features. The proposed method is also effective in analyzing mislocalization. For example, as shown in the first row of Fig. V, where the object detector mislocalizes the person by including the chair, the proposed method identifies both the features of the person and the towel covering the chair, indicating that these features contributed to the mislocalization. In contrast, other methods either highlight only one of these features or produce noisy explanations.
@@ D.13 Comparison with VPS
Figure W: Heat maps for an image containing multiple instances of the same category. VX-CODE heat maps are generated via patch insertion.
D.14 Heat maps comparison with multiple instances
A key difference between detector explanations and classification explanations is that the former should be instance-specific. To better demonstrate this property, we additionally provide qualitative results on images containing multiple objects from the same category. Specifically, we consider cases where the detector outputs several bounding boxes of the same class and visualize explanations for a selected detection.
Figure W shows the results. VX-CODE yields instance-focused evidence by capturing multiple complementary cues (e.g., the head and arms) that support the target detection. In contrast, several baselines tend to produce noisy or ambiguous explanations that are harder to interpret in such multi-instance scenes.
Figure X: Heat map comparisons for RetinaNet. VX-CODE heat maps are generated via patch insertion.
Table 12: Quantitative results on RetinaNet (100 detected instances from COCO). The best values are indicated in bold.
Method Ins $\uparrow$ Del $\downarrow$ OA $\uparrow$
Grad-CAM 0.890 0.420 0.470
Grad-CAM++ 0.899 0.407 0.492
DRISE 0.825 0.433 0.392
SSGrad-CAM 0.924 0.369 0.555
ODAM 0.929 0.354 0.575
SSGrad-CAM++ 0.932 0.348 0.584
VX-CODE 0.933 0.331 0.602
D.15 Results on a one-stage detector
In the main paper, we focus on two-stage, transformer-based detectors. To further verify that VX-CODE generalizes beyond this setting, we additionally evaluate it on the one-stage detector RetinaNet [31].
Figure X compares the heat map explanations produced by each method. VX-CODE also yields meaningful evidence regions on RetinaNet. Table 12 reports quantitative results computed on 100 detected instances from COCO. Overall, VX-CODE outperforms the baselines, demonstrating its effectiveness on a one-stage detector.
Table 13: Average pairwise interaction score II on COCO with DETR (300 instances). Higher values indicate stronger collective effects between selected patches.
Method Interaction score $\uparrow$
Grad-CAM 0.007
Grad-CAM++ 0.012
DRISE 0.031
SSGrad-CAM 0.017
ODAM 0.020
SSGrad-CAM++ 0.025
VX-CODE 0.051
D.16 Additional analysis of collective effects
We further analyze whether the selected patches exhibit collective effects. For two selected patches $i$ and $j$, we define the pairwise interaction score as
$I=f(\{i,j\})-f(\{i\})-f(\{j\})+f(\emptyset),$ (131)
where $f$ denotes the reward function. This corresponds to the interaction term in Def. 2 when $N=\{i,j\}$, and measures the synergy obtained by inserting patches $i$ and $j$ jointly into an empty image.
For each method, we aggregate the explanation heat map into patch-level scores, select the top-5 patches, and compute $I$ for all patch pairs. We evaluate this on COCO using DETR over 300 instances. Table 13 reports the average interaction score.
VX-CODE achieves the highest interaction score, indicating that it tends to select patches with stronger collective effects than the competing methods. This result complements the Del/Ins evaluation in the main paper and supports the effectiveness of VX-CODE in identifying jointly important evidence.